OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. Write your equation in slope-intercept form. Y – 10 = 3/2 (x – 4) ; (4, 2)

Nnesha (nnesha):

parallel lines have same slope

Nnesha (nnesha):

okay so first of all what is slope of this line Y – 10 = 3/2 (x – 4))???????

Nnesha (nnesha):

refresh your page if there is ???? question marks

OpenStudy (anonymous):

Y = 4x + 36

Nnesha (nnesha):

what is that ??

OpenStudy (anonymous):

That is the answer I got for this problem.

Nnesha (nnesha):

what is slope of the equation that yu have ??

Nnesha (nnesha):

like i said parallel lines have same slope so both equation should have same slope

OpenStudy (anonymous):

I messed up. The equation above should be: Y = 4x – 3; (-3, 12) That is what is in the question.

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

ohhh okay so what is slope of that equation

OpenStudy (anonymous):

I got Y = 4x + 24

OpenStudy (anonymous):

But the answer key says differently

Nnesha (nnesha):

yes it should be because htat is wrong :(

OpenStudy (anonymous):

What did I do wrong?

Nnesha (nnesha):

wait

Nnesha (nnesha):

yes your equation right

Nnesha (nnesha):

look at your question again ???

OpenStudy (anonymous):

Huh? @Nnesha

OpenStudy (anonymous):

Can you please explain to me how the answer would be Y = 4x + 24 ?

OpenStudy (aripotta):

wait this is the actual equation? Y = 4x – 3?

OpenStudy (anonymous):

The question is: Write an equation for the line that is parallel to the given line and passes through the given point. Write your equation in slope-intercept form. Y = 4x – 3; (-3, 12)

OpenStudy (aripotta):

ok so from looking at that equation, we know that our parallel line needs to have a slope of 4. since we know one point and the slope, we're going to put them in point-slope form, which is y - y1 = m(x - x1)

Nnesha (nnesha):

because parallel lines have same slope so 2nd equation should be like this y = 4x + b now plug x an dy values into the equation then solve for b

OpenStudy (anonymous):

What I don't understand is how the answer would be Y = 4x + 24

Nnesha (nnesha):